Solve the given differential equation by undetermined coefficients.
y'' − 2y' + 2y = e^2x(cos(x) − 8 sin(x))
Answer #1

y-2y! +2y = c2* (1052 - 8 Sima) Total solution gex) = get up 10 find yo, D2-20+2=0 then rie, m²- 2m +2=0. ma at 14-8 marbtJ61 COSX sim] you can be cosa] - 8e [ 8e2x [ Sina] ID²+20+2 LD?+20+2 COS(91+b) (a) Sincaut) fio put D²= -02 then 1 .-1+2D12 ba